OpenStudy (aryana_maria2323):

Please Help! Will FAN AND MEDAL! Use the graph of f(t) = 2t + 3 on the interval [−3, 6] to write the function F(x), where f of x equals the integral from 3 to x of f of t dt. a. F(x) = 2x2 + 6x b. F(x) = 2x + 3 c. F(x) = x2 + 3x + 54 d. F(x) = x2 + 3x − 18

OpenStudy (aryana_maria2323):

Can you please @jim_thompson5910 @agent0smith @ganeshie8

jimthompson5910 (jim_thompson5910):

`Use the graph of f(t) = 2t + 3 on the interval [−3, 6] to write the function F(x), where f of x equals the integral from 3 to x of f of t dt.` so \[\Large f(x) = \int_{3}^{x}f(t)dt\] \[\Large f(x) = \int_{3}^{x}(2t+3)dt\]

jimthompson5910 (jim_thompson5910):

You'll use the fundamental theorem of calculus \[\Large \int_{a}^{b}f(x)dx = g(b) - g(a)\] where g(x) = antiderivative of f(x)

OpenStudy (aryana_maria2323):

Okay so I have to plug in the second problem and find the anti-derivative

jimthompson5910 (jim_thompson5910):

what is the antiderivative of `2t+3` ?

OpenStudy (aryana_maria2323):

t^2+3t+C

jimthompson5910 (jim_thompson5910):

yep so let's make g(t) = t^2+3t+C

jimthompson5910 (jim_thompson5910):

now you need to compute g(x) and g(3) then subtract g(x) - g(3) = ??

OpenStudy (aryana_maria2323):

-3g

jimthompson5910 (jim_thompson5910):

g(t) = t^2+3t+C g(x) = x^2+3x+C ... replace every t with x

jimthompson5910 (jim_thompson5910):

g(t) = t^2+3t+C g(3) = (3)^2+3(3)+C ... replace every t with 3 g(3) = 9 + 9 + C g(3) = 18 + C

jimthompson5910 (jim_thompson5910):

so... g(x) - g(3) = [x^2+3x+C] - [18 + C] = ???

jimthompson5910 (jim_thompson5910):

tell me what you get or if you get stuck

OpenStudy (aryana_maria2323):

\[x^2+3x-18\] @jim_thompson5910

jimthompson5910 (jim_thompson5910):

correct. Nice work

OpenStudy (aryana_maria2323):

Can you help me with one more? @jim_thompson5910